# Moving Mesh with Streamline Upwind Petrov-Galerkin (MM-SUPG) Method for Time-dependent Convection-Dominated Convection-Diffusion Problems

### v3

@@ Abstract
Time-dependent convection-diffusion problems is considered, particularly when the diffusivity is very small and sharp layers exist in the solutions. Nonphysical oscillations may occur in the numerical solutions when using regular mesh with standard computational methods. In this work, we develop a moving mesh SUPG (MM-SUPG) method, which integrates the streamline upwind Petrov-Galerkin (SUPG) method with the moving mesh partial differential equation (MMPDE) approach. The proposed method is designed to handle both isotropic and anisotropic diffusivity tensors. For the isotropic case, we focus on improving the stability of the numerical solution by utilizing both artificial diffusion from SUPG and mesh adaptation from MMPDE. And for the anisotropic case, we focus on the positivity of the numerical solution. We introduce a weighted diffusion tensor and develop a new metric tensor to control the mesh movement. We also develop conditions for time step size so that the numerical solution satisfies the discrete maximum principle (DMP). Numerical results demonstrate that the proposed MM-SUPG method provides results better than SUPG with fixed mesh or moving mesh without SUPG.
@@ 1 Introduction
[…]
Here, $\mathbb{D}=\mathbb{D}({\boldsymbol{x}})$ on $\Omega_{T}=\Omega\times(0,T]$ is the diffusivity tensor that is symmetric and positive-definite (SPD), ${\boldsymbol{b}}$ represents an incompressible velocity field, while $g$, $u_{0}$, and $f$ are sufficiently smooth functions. The parameter $\varepsilon\in(0,1]$ is the scaling factor for the diffusivity coefficient. When $\mathbb{D}$ has the same eigenvalues, the problem is isotropic. In the anisotropic case, the eigenvalues of $\mathbb{D}$ are different, which indicates that diffusion in one direction is faster than in the other.
@@
When applying the finite difference method with only the center stencil, one obtains a numerical approximation that, depending on $\varepsilon$, tends to be under-diffusive. Conversely, utilizing only an upwind stencil leads to excessive diffusion, known as over-diffusive. As the exact solution typically lies somewhere in between these extremes, it prompts the adoption of upwinding the center scheme [7]. This method was initially proposed by Brooks and Hughes for incompressible Navier-Stokes problems [2], and has since proven to be fruitful in studying those equations [8, 9]. Comparison among different methods for the steady-state form of equation (1) have been performed in [10].
@@ 2.1 SUPG method
We consider an affine family of simplicial triangulations $\{\mathcal{T}_{h}\}$ for $\Omega$ and denote the piecewise linear finite element space as $V^{h}=(v\in H^{1}(\Omega),\,v|_{K}\in\mathcal{P}_{1}^{K},\forall K\in\mathcal{T}_{h})$, where $\mathcal{P}_{1}^{K}$ is the linear polynomial space in element $K$. Denote by $U^{h}=V^{h}\cap H_{g}^{1}(\Omega)$ and $U_{0}^{h}=V_{h}\cap H_{0}^{1}(\Omega)$. The standard Galerkin method is to find $u_{h}\in U^{h}$ such that
$\int_{\Omega}u_{t}v_{h}-\varepsilon\int_{\Omega}\nabla\cdot(\mathbb{D}\nabla u_{h})v_{h}+\sum_{K\in\mathcal{T}_{h}}({\boldsymbol{b}}\cdot\nabla u_{h})v_{h}=\int_{\Omega}fv_{h},\quad\forall v_{h}\in U_{0}^{h}.$ (4)
@@
where $\tau_{K}$ is the stabilization parameter. Therefore, one may perceive the space $U_{h}^{\textrm{SUPG}}$ as a suitable stabilized test space for singularly perturbed problems. The weak formulation for SUPG method is given as follows: find $u\in H^{1}(\Omega)$ such that
$\int_{\Omega}u_{t}v_{h}-\varepsilon\int_{\Omega}\nabla\cdot(\mathbb{D}\nabla u_{h})(v_{h}+\tau_{K}{\boldsymbol{b}}\cdot\nabla v_{h})+\sum_{K\in\mathcal{T}_{h}}({\boldsymbol{b}}\cdot\nabla u_{h})(v_{h}+\tau_{K}{\boldsymbol{b}}\cdot\nabla v_{h})=\int_{\Omega}f(v_{h}+\tau_{K}{\boldsymbol{b}}\cdot\nabla v_{h}),\quad\forall v_{h}\in U_{0}^{h}.$ (6)
A conventional choice of $\tau_{K}$ on mesh $\mathcal{T}_{h}$ is
@@
It determines local convection-dominated or diffusion-dominated regimes on the current mesh $\mathcal{T}_{K}$. There is flexibility in defining $\textrm{diam}(K)_{\mathcal{T}_{h}}$ and in selecting $c_{0}$ and $c_{1}$. This paper primarily focuses on scenarios where $\textrm{Pe}_{K}$ is large with respect to $\varepsilon^{-1}\|{\boldsymbol{b}}\|_{\infty}$. Typically, individual bilinear forms in the SUPG formulation are estimated, and $\tau_{\mathcal{T}_{h}}^{K}$ is derived based on those errors. In [3], such a method was employed using error estimates in the discrete norm
$\|w\|_{\varepsilon,k}^{2}=\varepsilon\|\nabla w\|^{2}+\sum_{K\in\mathcal{T}_{K}}\tau_{\mathcal{T}_{h}}^{K}\|{\boldsymbol{b}}\cdot\nabla w\|_{L^{2}(K)}^{2},\,\forall w\in H_{0}^{1}(\Omega)$ (9)
for time independent ${\boldsymbol{b}}$. However, our method relies on different error estimates for defining the metric tensor along with a mesh functional approach. If a source term is introduced, say $c\in L^{\infty}(0,T;L^{\infty}(\Omega))$, it has been proposed that the FE perturbation parameters should consider this source term [29, 30, 31, 32].
@@
$\tau_{\mathcal{T}_{h}}^{K}$ $=\frac{\textrm{diam}(K)_{\mathcal{T}_{K}}}{2\|{\boldsymbol{b}}\|_{\infty}}\xi(\textrm{Pe}_{K}),$ (10)
@@
$\textrm{diam}(K)=\sup_{{\boldsymbol{x}}_{1},{\boldsymbol{x}}_{2}\in K}\|{\boldsymbol{x}}_{1}-{\boldsymbol{x}}_{2}\|$ (11)
@@ 2.2 Moving mesh partial differential equations (MMPDE)
Moving mesh methods are instrumental in addressing domains where the solution exhibits excessively large gradient, such as sharp layers. Traditional methods like fitted operator methods, finite difference methods, finite element methods with uniform or Shishkin-type meshes often struggle to fully resolving these issues without a hyper-refined mesh or a priori knowledge about the solution. The MMPDE approach automatically concentrates mesh elements near the region where the solution exhibits large gradient, hence improves the computational efficiency. In this subsection, we give a brief overview of MMPDE. More details can be found in [33].
@@
$(\hat{K},\mathcal{P}_{1}^{\hat{K}},\Sigma_{\hat{K}})\xrightarrow{\quad F_{K}\quad}(K,\mathcal{P}_{1}^{K},\Sigma_{K}),$ (13)
For each element $K\in\mathcal{T}_{h}$, denote by $F_{K}:\hat{K}\to K$ the affine mapping between $K$ and the reference element $\hat{K}$ so that $\lvert\hat{K}\rvert=1$. Denote by ${\boldsymbol{x}}_{j}^{K}$ for $j=0$, 1, and 2 the vertices of $K$ and ${\boldsymbol{\xi}}_{j}^{K}$ the corresponding vertices of $\hat{K}$. Then the affine map $F_{K}$ maps ${\boldsymbol{\xi}}_{j}^{K}$ to ${\boldsymbol{x}}_{j}^{K}$, i.e. $F_{K}({\boldsymbol{\xi}}_{j}^{K})={\boldsymbol{x}}_{j}^{K}$ for $j=0$, 1, and 2. $\hat{K}$ is called the reference element and belongs to $\hat{\mathcal{T}}_{h}$, the reference mesh.
Denote by $\mathbb{M}_{K}$ the volume-average of $\mathbb{M}$ over element $K$, i.e
@@
    equidistribution: $\lvert K\rvert\sqrt{\det(\mathbb{M}_{K})}=\displaystyle\frac{\sigma_{h}}{N}$ for all $K\in\mathcal{T}_{h}$ and
@@
$\frac{\textrm{d}{\boldsymbol{x}}_{i}}{\textrm{d}t}=-\frac{P_{j}}{\gamma}\left(\frac{\partial I}{\partial{\boldsymbol{x}}_{i}}\right)^{T},\ i=1,\dots,N_{v},$ (16)
where $\gamma$ is a user parameter meant to adjust the time scale of the mesh movement and $P_{j}=\det\left(\mathbb{M}(x_{j})\right)^{\frac{p-1}{2}}$ is chosen so that the MMPDE is invariant under the scaling transformation of $\mathbb{M}$. The physical PDE and mesh PDE can be solved either simultaneously or alternatively. When the PDEs are solved simultaneously, the mesh can respond quickly to any change in the physical solution. However, this procedure is highly nonlinear. In our work, we employ the alternative approach. The mesh PDE is firstly solved based on the physical solution and the mesh at the current time step. Then mesh nodes are relocated based on the solutions from mesh PDE. Lastly, the physical PDE is solved on the new mesh for the next time step. The alternate procedure allows for flexibility and potential efficiency at each time level. One disadvantage of this approach is a lag in time, which can be addressed by employing multiple iterations on the same time level before advancing to the next. It is always beneficial to start with a initially adapted mesh, especially when sharp layers exist from the beginning.
@@
$\mathbb{M}=\det\left(\mathbb{I}+\lvert H_{K}(u)\rvert\right)^{-\frac{1}{6}}\left(\mathbb{I}+\lvert H_{K}(u)\rvert\right),$ (17)
@@
Then the polynomial $q$ in (Legendre) polynomial form
@@
Then $\nabla u({\boldsymbol{x}}_{j})\approx\nabla q({\boldsymbol{x}}_{j})$. This metric tensor is optimal in the $L^{2}$ norm of linear interpolation error on triangular meshes [35]. Moreover, it is clearly symmetric and one may replace $\lvert H_{K}(u)\rvert$ with $\alpha^{-1}|H_{K}(u)\rvert$, where $\alpha$ is called the intensity parameter and helps ensure that $\mathbb{M}$ is positive definite. This monitor function is derived on anisotropic error estimates (see [33], chapter 5, for a more detailed discussion and the derivation of (17) and choice of $\alpha$).
@@ 2.3 Moving mesh with Streamline Upwind Petrov-Galerkin (MM-SUPG) method
Now we apply the MMPDE method to solve the SUPG formulation (6). Write ${\boldsymbol{u}}_{h}({\boldsymbol{x}},t)=\sum\limits_{j=1}^{N}u_{j}(t)\varphi_{j}({\boldsymbol{x}})$, where $\{\varphi_{j}\}_{j=1}^{N}$ is the set of linear basis functions in $U_{h}^{0}$. Also, since $\varphi_{j}$ is a linear function, the element-wise integral average can be defined explicitly
$\frac{1}{|K|}\int_{K}\varphi_{j}\varphi_{i}\textrm{d}{\boldsymbol{x}}=\begin{cases}\displaystyle\frac{1}{(d+1)(d+2)}\textrm{\ \ for \ \ }i\neq j\\ \displaystyle\frac{2}{(d+1)(d+2)}\textrm{\ \ for \ \ }i=j\end{cases}.$ (21)
@@
The semi-discrete matrix formulation for (6) is given by
@@
where ${\boldsymbol{u}}=[u_{1},\dots,u_{Nv}]^{T}$, ${\boldsymbol{M}}$, ${\boldsymbol{A}}$, and ${\boldsymbol{b}}$ are given by
$M_{ij}$ $=\sum_{K\in\mathcal{T}_{h}}(\varphi_{j},\varphi_{i})_{0,K}+\tau_{K}(\varphi_{j},{\boldsymbol{b}}\cdot\nabla\varphi_{i})_{0,K}$ (26)
@@
$\quad\quad+\tau_{K}\left(({\boldsymbol{b}}\cdot\nabla\varphi_{j},{\boldsymbol{b}}\cdot\nabla\varphi_{i})_{0,K}+(\varepsilon\mathbb{D}_{K}\nabla\varphi_{j},\nabla({\boldsymbol{b}}\cdot\nabla\varphi_{i}))_{0,K}\right)$
@@
Denote $N_{t}$ as the number of time steps. We use a $\theta$ scheme to discretize in time. For (25), we approximate
${\boldsymbol{u}}_{h}^{\prime}(t)\approx\frac{{\boldsymbol{u}}_{h}(t_{m+1})-{\boldsymbol{u}}_{h}(t_{m})}{\Delta t},\ \ m=1,\dots,N_{t}.$ (29)
@@
$\tilde{{\boldsymbol{A}}}$ $={\boldsymbol{M}}+\theta\Delta t{\boldsymbol{A}}$ (33)
@@
In this work, we use $\theta=0.5$, which corresponds to the Crank-Nicholson. Fixed time step is used in the computations. In Section 2 we establish bounds on $\Delta t$ so that DMP is satisfied.

### v4

@@ Abstract
Time-dependent convection-dominated convection-diffusion problems are considered. We develop a moving mesh streamline upwind Petrov-Galerkin (MM-SUPG) method by combining residual-based SUPG stabilization with a metric-based moving mesh PDE (MMPDE) approach. The key feature of the method is the interaction between mesh adaptation and stabilization: the evolving mesh modifies both the direction and magnitude of the SUPG stabilization through the local element geometry, while stabilization improves robustness in convection-dominated regimes. For isotropic diffusion, numerical results show that the proposed method reduces spurious oscillations and provides improved resolution of sharp layers compared with fixed mesh methods, while yielding accuracy comparable to moving mesh finite element methods without SUPG in some cases. For anisotropic diffusion, we introduce a weighted tensor that incorporates both the diffusion tensor and the convection field, and construct a metric tensor via intersection to guide mesh adaptation. Under suitable assumptions on the elementwise tensors and mesh geometry, we establish sufficient conditions for the discrete maximum principle (DMP) of the fully discrete scheme. The analysis is based on quantitative bounds for the convection terms, and requires structural conditions on the alignment between the velocity field and the diffusion tensor. Numerical experiments demonstrate that the proposed metric improves monotonicity properties and reduces undershoots, while maintaining overall accuracy.
@@ 1 Introduction
$u_{t}-\varepsilon\nabla\cdot(\mathbb{D}\nabla u)+{\boldsymbol{b}}({\boldsymbol{x},t})\cdot\nabla u=f({\boldsymbol{x}},t)\textrm{\hskip 27.46295pt in \hskip 18.06749pt}\Omega\times(0,T],$ (1a)
$u({\boldsymbol{x}},t)=g({\boldsymbol{x}},t)\textrm{\hskip 130.08621pt on \hskip 18.06749pt}\partial\Omega\times[0,T],$ (1b)
$u({\boldsymbol{x}},0)=u_{0}({\boldsymbol{x}})\textrm{\hskip 133.69994pt in\hskip 18.06749pt}\Omega\times\{t=0\}.$ (1c)
Here, $\mathbb{D}=\mathbb{D}({\boldsymbol{x}})$ is the diffusivity tensor that is symmetric and positive-definite (SPD), $\boldsymbol{b}$} represents an incompressible velocity field, while $g$, $u_{0}$, and $f$ are sufficiently smooth functions. The parameter $\varepsilon\in(0,1]$ is the scaling factor for the diffusivity coefficient. When $\mathbb{D}$ has the same eigenvalues, the problem is isotropic. In the anisotropic case, the eigenvalues of $\mathbb{D}$ are different, which indicates that diffusion in one direction is faster than in the other.
@@
When applying the finite difference method with only the center stencil, one obtains a numerical approximation that, depending on $\varepsilon$, tends to be under-diffusive. Conversely, utilizing only an upwind stencil leads to excessive diffusion, known as over-diffusion. As the exact solution typically lies somewhere in between these extremes, it prompts the adoption of upwinding the center scheme [7]. This method was initially proposed by Brooks and Hughes for incompressible Navier-Stokes problems [2], and has since proven to be fruitful in studying those equations [8, 9]. Comparison among different methods for the steady-state form of equation (1) have been performed in [10].
@@
Recent developments include variational multiscale stabilized formulations combined with adaptive mesh strategies (e.g., [25]), as well as residual-based stabilization on anisotropic meshes. Compared with these approaches, the present work focuses on integrating SUPG stabilization with a metric-based moving mesh PDE framework, with particular emphasis on discrete maximum principle (DMP) preservation for anisotropic diffusion.
@@
The novelty of the proposed MM-SUPG method lies not merely in combining moving mesh adaptation with SUPG stabilization, but in analyzing and exploiting their interaction. The effectiveness of SUPG stabilization depends strongly on mesh geometry, as both the streamline derivative ${\boldsymbol{b}}\cdot\nabla u_{h}$ and the stabilization parameter $\tau_{K}$ are mesh-dependent. Meanwhile, the moving mesh method dynamically aligns elements according to solution features and anisotropic diffusion. This coupling leads to a nontrivial interplay: mesh alignment influences the direction and strength of SUPG stabilization, which in turn affects oscillation control and monotonicity. In particular, for anisotropic diffusion, the proposed metric tensor construction balances alignment with both the convection field and diffusion tensor, enabling improved discrete maximum principle behavior. This interaction between mesh adaptation and stabilization is a central contribution of the present work.
@@ 2.1 SUPG method
We consider an affine family of simplicial triangulations $\{\mathcal{T}_{h}\}$ for $\Omega$ and denote the piecewise linear finite element space as $V^{h}=(v\in H^{1}(\Omega),\,v|_{K}\in\mathcal{P}_{1}^{K},\forall K\in\mathcal{T}_{h})$, where $\mathcal{P}_{1}^{K}$ is the linear polynomial space in element $K$. Denote by $U^{h}=V^{h}\cap H_{g}^{1}(\Omega)$ and $U_{0}^{h}=V^{h}\cap H_{0}^{1}(\Omega)$. The standard Galerkin method is to find $u_{h}\in U^{h}$ such that
$\int_{\Omega}(u_{h})_{t}v_{h}-\varepsilon\sum_{K\in\mathcal{T}_{h}}\nabla\cdot(\mathbb{D}_{K}\nabla u_{h})v_{h}+\int_{\Omega}({\boldsymbol{b}}\cdot\nabla u_{h})v_{h}=\int_{\Omega}fv_{h},\quad\forall v_{h}\in U_{0}^{h},$ (4)
where $\mathbb{D}_{K}$ is the restriction of $\mathbb{D}$ to element $K$.
@@
where $\tau_{K}$ is the stabilization parameter. Therefore, one may perceive the space $U_{h}^{\textrm{SUPG}}$ as a suitable stabilized test space for singularly perturbed problems. The weak formulation for SUPG method is given as follows: find $u_{h}\in U^{h}(\Omega)$ such that
$\int_{\Omega}(u_{h})_{t}v_{h}$ $-\varepsilon\sum_{K\in\mathcal{T}_{h}}\nabla\cdot(\mathbb{D}_{K}\nabla u_{h})(v_{h}+\tau_{K}{\boldsymbol{b}}\cdot\nabla v_{h})$
$+\sum_{K\in\mathcal{T}_{h}}({\boldsymbol{b}}\cdot\nabla u_{h})(v_{h}+\tau_{K}{\boldsymbol{b}}\cdot\nabla v_{h})=\int_{\Omega}f(v_{h}+\tau_{K}{\boldsymbol{b}}\cdot\nabla v_{h}),\quad\forall v_{h}\in U_{0}^{h}.$ (6)
The stabilization parameter $\tau_{K}$ is chosen following classical SUPG formulations, where it scales with the local element size and inverse flow magnitude to balance numerical diffusion and consistency; see [2, 8, 30], and related developments in moving mesh contexts [3]. A conventional choice of $\tau_{K}$ on mesh $\mathcal{T}_{h}$ is
@@
$\tau_{K}$ determines local convection-dominated or diffusion-dominated regimes on the current mesh $\mathcal{T}_{K}$. There is flexibility in defining $\textrm{diam}(K)_{\mathcal{T}_{h}}$ and in selecting $c_{0}$ and $c_{1}$. This paper focuses on convection-dominated regimes characterized by large element Péclet numbers $P_{e_{K}}\gg 1$, which may arise due to small $\varepsilon$ and/or insufficient mesh resolution relative to the flow magnitude.. Typically, individual bilinear forms in the SUPG formulation are estimated, and $\tau_{\mathcal{T}_{h}}^{K}$ is derived based on those errors. In [3], such a method was employed using error estimates in the discrete norm for time independent $\boldsymbol{b}$}
$\|w\|_{\varepsilon,k}^{2}=\varepsilon\|\nabla w\|^{2}+\sum_{K\in\mathcal{T}_{K}}\tau_{\mathcal{T}_{h}}^{K}\|{\boldsymbol{b}}\cdot\nabla w\|_{L^{2}(K)}^{2},\,\forall w\in H_{0}^{1}(\Omega).$ (9)
However, our method relies on different error estimates for defining the metric tensor along with a mesh functional approach. When a reaction/source term is present, stabilization parameters must be modified to account for its contribution to the residual, as discussed in [31, 32, 33, 34].
@@
$\tau_{\mathcal{T}_{h}}^{K}$ $=\frac{\textrm{diam}(K)_{\mathcal{T}_{h}}}{2\|{\boldsymbol{b}}\|_{\infty}}\xi(\textrm{Pe}_{K}),$ (10)
@@
$\textrm{diam}(K)=\sup_{{\boldsymbol{x}}_{1},{\boldsymbol{x}}_{2}\in K}\|{\boldsymbol{x}}_{1}-{\boldsymbol{x}}_{2}\|,$ (11)
@@ 2.2 Moving mesh partial differential equations (MMPDE)
Moving mesh methods are instrumental in addressing domains where the solution exhibits excessively large gradient, such as sharp layers. Traditional methods like fitted operator methods, finite difference methods, finite element methods with uniform or Shishkin-type meshes may not be able to catch the sharp gradient without a hyper-refined mesh or a priori knowledge about the solution. The MMPDE approach automatically concentrates mesh elements near the region where the solution exhibits large gradient, hence improves the computational efficiency. In this subsection, we give a brief overview of MMPDE. More details can be found in [35].
@@
$(\hat{K},\mathcal{P}_{1}^{\hat{K}},\Sigma_{\hat{K}})\xrightarrow{\quad F_{K}\quad}(K,\mathcal{P}_{1}^{K},\Sigma_{K}).$ (13)
For each element $K\in\mathcal{T}_{h}$, denote by $F_{K}:\hat{K}\to K$ the affine mapping between $K$ and the reference element $\hat{K}$ so that $\lvert\hat{K}\rvert=1$. Denote by ${\boldsymbol{x}}_{j}^{K}$ for $j=0$, 1, and 2 the vertices of $K$ and ${\boldsymbol{\xi}}_{j}^{K}$ the corresponding vertices of $\hat{K}$. Then the affine map $F_{K}$ maps ${\boldsymbol{\xi}}_{j}^{K}$ to ${\boldsymbol{x}}_{j}^{K}$, i.e., $F_{K}({\boldsymbol{\xi}}_{j}^{K})={\boldsymbol{x}}_{j}^{K}$ for $j=0$, 1, and 2. $\hat{K}$ is called the reference element and belongs to $\hat{\mathcal{T}}_{h}$, the reference mesh.
Denote by $\mathbb{M}_{K}$ the volume-average of $\mathbb{M}$ over element $K$, i.e.,
@@
    equidistribution: $\lvert K\rvert\sqrt{\det(\mathbb{M}_{K})}=\displaystyle\frac{\sigma_{h}}{N}$ for all $K\in\mathcal{T}_{h}$, and
@@
$\frac{\textrm{d}{\boldsymbol{x}}_{i}}{\textrm{d}t}=-\frac{P_{i}}{\gamma}\left(\frac{\partial I}{\partial{\boldsymbol{x}}_{i}}\right)^{T},\ i=1,\dots,N_{v},$ (16)
where $\gamma$ is a user parameter meant to adjust the time scale of the mesh movement and $P_{j}=\det\left(\mathbb{M}(x_{j})\right)^{\frac{p-1}{2}}$ is chosen so that the MMPDE is invariant under the scaling transformation of $\mathbb{M}$.
The physical PDE and mesh PDE can be solved either simultaneously or alternatively. When the PDEs are solved simultaneously, the mesh can respond quickly to any change in the physical solution. However, this procedure is highly nonlinear. In our work, we employ the alternative approach. The mesh PDE is firstly solved based on the physical solution and the mesh at the current time step. Then mesh nodes are relocated based on the solutions from mesh PDE. Lastly, the physical PDE is solved on the new mesh for the next time step. The alternate procedure allows for flexibility and potential efficiency at each time level. One disadvantage of this approach is a lag in time, which can be addressed by employing multiple iterations on the same time level before advancing to the next. It is always beneficial to start with a initially adapted mesh, especially when sharp layers exist from the beginning.
@@
[…]
@@
Then the polynomial $q$ is in (Legendre) polynomial form
@@
Then it is clear that $\nabla u({\boldsymbol{x}}_{j})\approx\nabla q({\boldsymbol{x}}_{j})$. This metric tensor is optimal in the $L^{2}$ norm of linear interpolation error on triangular meshes [37]. Moreover, it is clearly symmetric and one may replace $\lvert H_{K}(u)\rvert$ with $\alpha^{-1}|H_{K}(u)\rvert$, where $\alpha$ is called the intensity parameter and helps ensure that $\mathbb{M}$ is positive definite. This monitor function is derived on anisotropic error estimates (see [35], chapter 5, for a more detailed discussion and the derivation of (17) and choice of $\alpha$).
@@ 2.3 Moving mesh with Streamline Upwind Petrov-Galerkin (MM-SUPG) method
Now we apply the MMPDE method to solve the SUPG formulation (2.1). Write ${\boldsymbol{u}}_{h}({\boldsymbol{x}},t)=\sum\limits_{j=1}^{N_{v}}u_{j}(t)\varphi_{j}({\boldsymbol{x}})$, where $\{\varphi_{j}\}_{j=1}^{N_{v}}$ is the set of linear basis functions in $U_{0}^{h}$. Also, since $\varphi_{j}$ is a linear function, the element-wise integral average can be defined explicitly
[…]
@@
The semi-discrete matrix formulation for (2.1) is given by
@@
where ${\boldsymbol{u}}=[u_{1},\dots,u_{N_{v}}]^{T}$, $\boldsymbol{M}$}, $\boldsymbol{A}$}, and $\boldsymbol{f}$} are given by
$M_{ij}$ $=\sum_{K\in\mathcal{T}_{h}}(\varphi_{j},\varphi_{i})_{0,K}+\tau_{K}(\varphi_{j},{\boldsymbol{b}}\cdot\nabla\varphi_{i})_{0,K},$ (26)
@@
$\quad\quad+\tau_{K}\left(({\boldsymbol{b}}\cdot\nabla\varphi_{j},{\boldsymbol{b}}\cdot\nabla\varphi_{i})_{0,K}+(\varepsilon\mathbb{D}_{K}\nabla\varphi_{j},\nabla({\boldsymbol{b}}\cdot\nabla\varphi_{i}))_{0,K}\right),$
@@
We use a $\theta$-scheme to discretize in time. For (25), we approximate
${\boldsymbol{u}}_{h}^{\prime}(t_{m})\approx\frac{{\boldsymbol{u}}_{h}(t_{m+1})-{\boldsymbol{u}}_{h}(t_{m})}{\Delta t},\ \ m=1,\dots,N_{t}.$ (29)
@@
$\tilde{{\boldsymbol{A}}}$ $={\boldsymbol{M}}+\theta\Delta t{\boldsymbol{A}},$ (33)
@@
In this work, we use $\theta=0.5$, which corresponds to the Crank-Nicolson method. Fixed time step is used in the computations. In Section 4 we derive conditions on $\Delta t$ so that DMP is satisfied.
To generate the adapted mesh at each physical time level, we minimize the discrete meshing functional $I(\mathcal{T}_{h})$ defined in (14) by solving the mesh PDE (16). The derivatives of the meshing functional are assembled element-by-element following the MMPDE framework in [35, 37] and the mesh PDE can be written in the form
$\frac{\textrm{d}{\boldsymbol{x}}_{i}}{\textrm{d}t}=\frac{P_{i}}{\gamma}\sum_{K\in\omega_{i}}|K|{\boldsymbol{v}}_{i_{K}}^{K},\ i=1,\dots,N_{v},$ (35)
where $\omega_{i}$ is the element patch associated with the $i$-th vertex, $i_{K}$ is the local index of vertex $i$ in element $K$, and ${\boldsymbol{v}}_{i_{K}}^{K}$ denotes the local mesh velocity contributed by element $K$. Here the local velocities are obtained from the derivatives of $G(\mathbb{J}_{K},\det(\mathbb{J}_{K}),\mathbb{M}_{K})$ defined in (15) with respect to $\mathbb{J}_{K}$ and $\det(\mathbb{J}_{K})$, where $\mathbb{J}_{K}=(F_{K}^{\prime})^{-1}$. Boundary vertices are constrained to remain on $\partial\Omega$, thus boundary nodes are allowed to slide along the boundary with corner nodes being fixed.
To obtain the new mesh, we integrate the discretized mesh PDE (35) from the current mesh $\mathcal{T}_{h}^{m}$ to the updated mesh $\mathcal{T}_{h}^{m+1}$. In our implementation, this system of ODEs is solved in pseudo-time using MATLAB ode15s solver. After the integration is completed, the new mesh $\mathcal{T}_{h}^{m+1}$ is obtained directly from the updated vertex coordinates.
A key point is that the matrices $\boldsymbol{M}$} and $\boldsymbol{A}$}, and the stabilization parameter $\tau_{K}$ are all assembled on the current physical mesh. Since the mesh changes in time, these quantities are recomputed after each mesh update. In particular, $\tau_{K}$ is evaluated on the updated mesh, so the amount of streamline stabilization adapts dynamically to the local mesh geometry. We emphasize that we do not solve the MMPDE and the physical PDE fully simultaneously. Instead, we use a staggered procedure:
(1) fix the physical solution and update the mesh;
(2) freeze the new mesh and solve the SUPG system for the physical solution.
This decouples the nonlinear mesh movement from the stabilized finite element solution and is considerably simpler and more robust in practice. The algorithm is given as follows.
Algorithm 1: MM-SUPG Method
for $m=1,...,N_{t}$, do:
$t_{m+1}=t_{m}+\Delta t$,
  • (1)
    Metric construction.
    Using ${\boldsymbol{u}}_{h}^{m}$ on $\mathcal{T}_{h}^{m}$, construct the metric tensor $\mathbb{M}^{m}=\mathbb{M}({\boldsymbol{u}}_{h}^{n})$.
  • (2)
    Mesh movement.
    Fixing $\mathbb{M}^{m}$, solve the MMPDE for the mesh coordinates from ${\boldsymbol{x}}^{m}$ to an updated vertex set ${\boldsymbol{x}}^{m+1}$, yielding the new mesh $\mathcal{T}_{h}^{m+1}$.
  • (3)
    Solution interpolation.
    Interpolate ${\boldsymbol{u}}_{h}^{m}$ from $\mathcal{T}_{h}^{m}$ onto $\mathcal{T}_{h}^{m+1}$, denoted by $\tilde{{\boldsymbol{u}}}_{h}^{m}$.
  • (4)
    Recompute geometric quantities.
    Using $\tilde{{\boldsymbol{u}}}_{h}^{m}$ on $\mathcal{T}_{h}^{m+1}$, compute $\textrm{diam}(K)_{\mathcal{T}_{h}}$, $\tau_{\mathcal{T}_{h}}^{K}$ and basis-function gradients.
  • (5)
    Assemble SUPG system.
    Assemble the mesh-dependent matrices ${\boldsymbol{M}}^{M+1}$, ${\boldsymbol{A}}^{m+1}$, and ${\boldsymbol{f}}^{m+1}$.
  • (6)
    Solve for physical solution.
    Solve the following linear system for the solution at $t_{m+1}$.
    ${\boldsymbol{M}}^{m+1}+\theta\Delta t{\boldsymbol{A}}^{m+1}{\boldsymbol{u}}_{h}^{m+1}=\theta\Delta t{\boldsymbol{f}}^{m+1}+\Delta t(1-\theta){\boldsymbol{f}}^{m}+\left({\boldsymbol{M}}^{m+1}-\Delta t(1-\theta){\boldsymbol{A}}^{m+1}\right){\boldsymbol{u}}_{h}^{m}.$
The effectiveness of MM-SUPG stems from the complementary roles of its components: moving mesh adaptation reduces interpolation error by aligning elements with solution features, while SUPG stabilization suppresses spurious oscillations along streamlines. Together, they yield improved stability without excessive numerical diffusion.